OpenStudy (anonymous):

Integrate 1/sqrt(3-x) from 2 to 3

myininaya (myininaya):

\[\int\limits_{2}^{3}\frac{1}{\sqrt{3-x}} dx\]

myininaya (myininaya):

let u=3-x => du=-1 dx

myininaya (myininaya):

so if du=-1 dx then -du=1 dx

myininaya (myininaya):

\[\int\limits_{3-2}^{3-3}\frac{-du}{\sqrt{u}}\]

myininaya (myininaya):

\[-\int\limits_{1}^{0}u^\frac{-1}{2} du=-2u^\frac{1}{2}|_1^0=-2(0^\frac{1}{2}-1^\frac{1}{2})=-2(-1)=2\]